# Investigating Efficacy of Perplexity in Detecting LLM-Generated Code

### v1

@@
Abstract
Large language model-generated code (LLMgCode) has become increasingly prevalent in software development. Many studies report that LLMgCode has more quality and security issues than human-authored code (HaCode). It is common for LLMgCode to mix with HaCode in a code change, while the change is signed by only human developers, without being carefully checked. Many automated methods have been proposed to detect LLMgCode from HaCode, in which the perplexity-based method (Perplexity for short) is the state-of-the-art method. However, the efficacy evaluation of Perplexity has focused on the detection accuracy. In this article, we are interested in whether Perplexity is good enough in a wider range of realistic evaluation settings. To this end, we devise a large-scale dataset that includes 11,664 HaCode snippets and 13,164 LLMgCode snippets, and based on that, we carry out a family of experiments to compare Perplexity against feature-based and pre-training-based methods from three perspectives: (1) detection accuracy in terms of programming language, degree of difficulty, and scale of solution, (2) generalization capability, and (3) inference efficiency. The experimental results show that Perplexity has the best generalization capability while it has low accuracy and efficiency in most cases. Based on the experimental results and detection mechanism of Perplexity, we discuss implications into both the strengths and limitations of Perplexity, e.g., Perplexity is unsuitable for high-level programming languages while it has good interpretability. As the first large-scale investigation on detecting LLMgCode from HaCode, this article provides a wide range of evidence for future improvement.
Index Terms:
@@ Abstract
I Introduction
@@
Large language model-generated code (LLMgCode) has become increasingly common in daily software development [1, 2]. Although LLMgCode reduce manual coding efforts and help developers focus on high-level tasks [3], the quality and security of LLMgCode are challenging [4, 5, 6, 7]. According to Pearce et al. [4], 40.73% of the 1,689 programs generated by GitHub Copilot are vulnerable, with risks far more than human-authored code (HaCode) [8, 9]. Moreover, LLMgCode has the risks of plagiarism and copyright infringement [10], which are less in HaCode. According to Guo et al. [11], LLMgCode can be a copy of HaCode even if the latter is with copyright protection [12]. Due to this, GitHub Copilot is being accused of compensating \$9 billion for generating restrictive code without attaching the necessary copyright or licenses [13].
@@
In addition to the challenges of LLMgCode itself, the mixture of LLMgCode and HaCode has two challenges. The first is compromising the transparency of software development, which results in inaccurate attribution of code ownership in bug accountability and obscured workload assessment in creating software artifacts [14]. The second is increasing the difficulty of evaluating the reliability of third-party software, which results in distrust for upstream dependencies in the software supply chain [15]. As the authorship of LLMgCode has not been provided, detecting LLMgCode can be a stop-gap [15]. To this end, many studies [14, 16] and commercial tools [17] have proposed to detect LLMgCode from LLMgCode, which supports mitigating various risks by taking measures such as conducting stricter security inspections and performing code plagiarism detection [18].
@@
The past few years have seen many studies [14, 15, 19, 20, 21, 22, 23, 24, 25] that proposed methods to detect LLMgCode from HaCode. According to the detection mechanisms used, these methods can be divided into three categories: 1. feature-based, 2. pre-training-based, and 3. perplexity-based methods (Perplexity for short for the last category). Feature-based methods [15, 19] are generally conventional machine learning classifiers that require professionals to devise appropriate features from aspects such as lexical and syntactic characteristics. Pre-training-based methods [20] take the code snippets as inputs. They convert code snippets to tokens, embed them as vectors, and use the pre-trained model as a classifier. Perplexity  [^1] For succinctness, Perplexity refers to the family of perplexity-based methods throughout the paper unless it is specifically stated otherwise. [14, 21, 22, 23, 24, 25] is a class of heuristic methods, which assume that LLMgCode has a lower perplexity than HaCode. Technically, perplexity is a measure of how confident a probability model predicts a sample [26]. When it comes to code generation, perplexity is used to measure the predictability of the next token of code based on prompts and existing tokens. The calculation of the perplexity is based on an LLM.
Figure 1: An example of calculating the perplexity of existing code
@@
Figure 1 presents an example of calculating the perplexity of a probability model (CodeLlama) for code snippets (either LLMgCode or HaCode). The prompt used is shown in lines 2-3, and the code snippets to be calculated are in line 6. Based on the prompt, LLM calculates the probability of “  ” (Tab) as the first token is 0.20. Based on the prompt and “  ”, the LLM calculates the probability of “return” as the second token is 0.31. Followed by “ tf” and “.train” are 0.98 and 0.99, respectively. The higher the probability of a token, the more confident the model is about generating this token. Therefore, the perplexity of the model in understanding this token is low. Perplexity is a probability-based measure and has many implementations (e.g., log of probability [21], log of rank [23], and log of probability divided by log of rank [25]) in which the log of probability is common. For example, the perplexity of CodeLlama for “return” is $ln(0.31)=-1.18$. The perplexity of a model for a piece of code is the average perplexity of tokens.
@@
When the code is natural and coherent, the model is confident in predicting the next token based on the prompts and the existing tokens, resulting in a lower perplexity. According to the mechanism of LLMs for code generation, LLMs generate code in the unit of a token through understanding previous tokens, calculating the probability of each possible next token, and applying the sampling strategy to select one with high probability as the output. Compared with LLMgCode, HaCode is more casual and personalized, making it more unpredictable. Therefore, it is reasonable that LLMgCode has lower perplexity than HaCode. Due to its explainability and scalability (e.g., log of probability, log of rank), Perplexity has been widely used [21, 22, 23, 24, 25].
@@
Although Perplexity is becoming the most popular method, the efficacy evaluation of Perplexity is limited to detection accuracy. For example, Shi et al. [14] evaluated their proposed DetectCodeGPT using AUC (Area Under the Curve) on Python only. Yet it is unclear the accuracy (effectiveness) of DetectCodeGPT other than Python. DetectCodeGPT detects only the code generated by LLMs with fewer than 7B parameters, its effectiveness on the commonly used commercial models such as GPT-4o and Gemini is unclear. Moreover, DetectCodeGPT uses the same model for code generation and perplexity calculation. As it is hard to early access models for code generation, the effectiveness of DetectCodeGPT in a wide range of contexts is unclear. The limited evaluation of Perplexity does not reflect its effects well in practice.
@@
Toward realistic evaluations of Perplexity, we conduct a large-scale investigation on Perplexity from three perspectives: detection accuracy, generalization capability, and inference efficiency. Detection accuracy is the foundation of detecting LLMgCode, while it fluctuates between situations, e.g., an effective method for C++ may not suit Python. Therefore, we evaluate the accuracy in various situations from three aspects. Specifically, the detection accuracy (effectiveness) is evaluated from the aspects of the programming language (C, C#, C++, Go, Java, Python, Ruby, and Rust), degree of difficulty (easy, middle, hard), and scale of solution (small, middle, and large). As new powerful LLMs are emerging, developers utilize various LLMs to assist in programming, therefore the generalization capability of Perplexity is important. The generalization capability is evaluated from the aspect of code generation models (GPT-3.5, GPT-4o, Gemini, and Llama). The waiting time greatly affects the user experience, thereby the inference efficiency of Perplexity should be considered. The inference efficiency is evaluated from the aspect of time spent (seconds/sample). We hope the multi-perspective efficacy evaluations can reflect the effects of Perplexity in practice. With experimental evaluation, we first devised a dataset containing 729 programming puzzles. Then, we collected 11,664 HaCode solutions from CodeNet [27] and generated 11,664 LLMgCode solutions using GPT-3.5. Next, we compared eight perplexity-based methods against two feature-based methods and one pre-training-based method that have been used to detect LLMgCode from HaCode.
@@
The experimental results show that (1) for effectiveness, the best-performing Perplexity performs well for detecting LLMgCode in C/C++ (AUC=91%) and large-scale code (AUC=88%) but still does not perform as well as the other two methods regarding programming language, degree of difficulty, and scale of solution; (2) for generalization capability, Perplexity perform better than others across different LLMs, with AUC ranging from 80% to 82%; (3) for inference efficiency, Perplexity without perturbation is as efficient as others, while Perplexity with perturbation have a relatively slow efficiency, with detection speed ranging from 10 to 150 seconds per sample. Based on the experimental results, we provide some implications of Perplexity from both positive and negative perspectives. On the negative side, Perplexity is not suitable for high-level programming languages, short code snippets, and just-in-time detection. On the positive side, Perplexity has good interpretability, exhibits good generalization capability regarding LLMs, and has the potential for fine-grained detection.
@@
  • Experimental Design: An experimental design to investigate the detection accuracy, generation capability, and inference efficiency of Perplexity for detecting LLMgCode.
  • Evidence: A wide range of evidence that illustrates advantages and limitations of Perplexity for detecting LLMgCode.
@@
The remainder of this article is organized as follows. Section II presents background and related work. Section III illustrates motivating examples. Section IV and Section V elaborate on the research methodology and results, respectively. Section VI discusses implications. Finally, we present limitations in Section VII and conclude this article in Section VIII.
II Background and Related Work
@@
This section gives an overview of LLMs for code generation and reviews studies that apply Perplexity for code authorship attribution.
II-A Background: Large Language Models for Code Generation
@@
Since LLMs were proposed, they have been widely applied in many tasks across a variety of domains [28, 29, 30, 31]. One of the most exciting applications of LLMs is code generation [32, 33, 34, 35, 36]. LLMs are capable of generating source code based on the prompts. The prompts consist of natural and programmatic descriptions of code functions, retractions, input and output formats, or any other elements. Following the prompts, an LLM generates tokens sequentially until it arrives at a stop indicator or arrives at the token limit.
@@
Many studies show that LLMgCode helps save time [37], reduce effort [38], and improve productivity [39] in software development. Moreover, LLMgCode help ease developers as they sometimes suffer from heavy workload [40]. With LLMgCode, developers can pay more attention to important and urgent tasks. Because of these benefits, LLMgCode has been widely used and even become a must-have in modern software development [41, 42]. A survey made with 500 USA developers showed that 92% of respondents use LLMgCode for both official assignments and personal use [43].
@@
Although LLMs bring a lot of benefits, the quality and security of LLMgCode remain challenging [4, 44, 45]. One of the causes is the corpus used for training LLMs is not qualified and secure enough [5]. Therefore, LLMgCode might be buggy and vulnerable [46, 47]. For example, Fu et al. [48] assessed the security of LLMgCode in six common programming languages and found that 156 out of 435 code segments contained security vulnerabilities. The proportion of vulnerabilities in C++ and Go code accounts for 46.1% and 45.0%, respectively. Mastropaolo et al. [49] assessed the robustness of LLMgCode and found that slight modifications to the natural language description while still maintaining semantic equivalence would result in the LLMs generating different results, occurring in 46% (408/892) of the total samples. Jimenez et al. [50] assessed the capability of large language models to solve 2294 real-world problems (Issues) on GitHub and found that the pass rate of code generated using GPT-4 has the lowest pass rate–0%, using Claude 2 has the highest pass rate–1.96%. In addition, LLMgCode might be similar or even unchanged to existing code snippets in the training corpus [11]. As the training corpus might include a significant amount of restrictive code [12], there are copyright infringement risks for LLMgCode.
II-B Related Work: Perplexity for Code Authorship Attribution

### v2

# One Size Does Not Fit All: Investigating Efficacy of Perplexity in Detecting LLM-Generated Code

@@
Abstract.
Large language model-generated code (LLMgCode) has become increasingly common in software development. So far LLMgCode has more quality issues than human-authored code (HaCode). It is common for LLMgCode to mix with HaCode in a code change, while the change is signed by only human developers, without being carefully examined. Many automated methods have been proposed to detect LLMgCode from HaCode, in which the perplexity-based method (Perplexity for short) is the state-of-the-art method. However, the efficacy evaluation of Perplexity has focused on detection accuracy. Yet it is unclear whether Perplexity is good enough in a wider range of realistic evaluation settings. To this end, we carry out a family of experiments to compare Perplexity against feature- and pre-training-based methods from three perspectives: detection accuracy, detection speed, and generalization capability. The experimental results show that Perplexity has the best generalization capability while having limited detection accuracy and detection speed. Based on that, we discuss the strengths and limitations of Perplexity, e.g., Perplexity is unsuitable for high-level programming languages. Finally, we provide recommendations to improve Perplexity and apply it in practice. As the first large-scale investigation on detecting LLMgCode from HaCode, this article provides a wide range of findings for future improvement.
@@ Abstract.
journal: JACMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Software and its engineering Software verification and validationccs: Software and its engineering Software supply chain security
1. Introduction
The advent of Large Language Models (LLMs) has profoundly advanced automated software engineering (Chen et al., 2021), of which code generation is most popular (Gu et al., 2024). Code generation is to generate source code based on natural descriptions (Li et al., 2023). Various LLMs (e.g., GitHub Copilot (Github, 2025), GPT-4 (OpenAI, 2025)) have been proposed for code generation in recent years (Du et al., 2024). LLMs have the strong capability to follow instructions (Chung et al., 2024), enabling developers to write code by providing natural language descriptions of requirements only (Jiang et al., 2024). Leveraging natural language understanding and generative power, LLMs achieve remarkable success in code generation (Austin et al., 2021). Therefore, LLM-generated code (LLMgCode) has become increasingly common in software development (Sun et al., 2024; Yu et al., 2024a).
@@ 1. Introduction
Although LLMgCode help developers reduce manual efforts and focus on high-level tasks (Liu et al., 2024a), the quality and security of LLMgCode are challenging (Pearce et al., 2022; Asare et al., 2023; Fu et al., 2023; Liu et al., 2024b). According to Pearce et al. (2022), 40.73% of the 1,689 programs generated by GitHub Copilot are vulnerable, with risks far more than human-authored code (HaCode(Perry et al., 2024; Sandoval et al., 2023). Moreover, LLMgCode has the risks of plagiarism and copyright infringement (Guo et al., 2023), which are less in HaCode. According to Guo et al. (2024), LLMgCode can be a copy of HaCode even if the latter is with copyright protection (Yu et al., 2023). Due to this, GitHub Copilot was accused of compensating \$9 billion for generating restrictive code without attaching the necessary copyright or licenses (Hale, 2022).
@@
In addition to the challenges of LLMgCode itself, the mixture of LLMgCode and HaCode has two challenges. The first is compromising the transparency of software development, which results in inaccurate attribution of code ownership in bug accountability and obscured productivity assessments (Shi et al., 2025). The second is increasing the difficulty of evaluating the reliability of third-party software, which results in distrust for upstream dependencies in the software supply chain (Bukhari et al., 2023). Therefore, It is important to detect LLMgCode from HaCode. To this end, many studies (Shi et al., 2025; Suh et al., 2025) and commercial tools (Copyleaks, 2025) have proposed to detect LLMgCode from HaCode, which supports mitigating various risks by taking measures such as conducting stricter security inspections and performing code plagiarism detection (Xu et al., 2025).
@@
The past few years have seen many studies (Shi et al., 2025; Bukhari et al., 2023; Idialu et al., 2024; Nguyen et al., 2024; Solaiman et al., 2019; Gehrmann et al., 2019; Ippolito et al., 2019; Mitchell et al., 2023; Su et al., 2023) that proposed methods to detect LLMgCode from HaCode. According to the detection mechanisms these methods use, they can be divided into three categories: (1) feature-based, (2) pre-training-based, and (3) perplexity-based methods (Perplexity for short). Feature-based methods (Bukhari et al., 2023; Idialu et al., 2024) are generally conventional machine learning classifiers that require professionals to devise appropriate features from various aspects such as lexical and syntactic characteristics. Pre-training-based methods (Nguyen et al., 2024) take the code snippets as inputs. They convert code snippets to tokens, embed them as vectors, and use the pre-trained model as a classifier. Perplexity  [^1] For succinctness, Perplexity refers to the family of perplexity-based methods throughout the paper unless it is specifically stated otherwise. (Shi et al., 2025; Solaiman et al., 2019; Gehrmann et al., 2019; Ippolito et al., 2019; Mitchell et al., 2023; Su et al., 2023) is a class of heuristic methods, which assume that LLMgCode has a lower perplexity than HaCode. Technically, perplexity is a measure of how confident a probability model is in predicting a sample (Xu and Sheng, 2024). When it comes to code generation, perplexity is used to measure the predictability of the next token of code based on prompts and existing tokens. The calculation of the perplexity is based on an LLM.
Figure 1. An example of calculating the perplexity of a set of existing code snippets
@@
Figure 1 presents an example of calculating the perplexity of a probability model (CodeLlama) for code snippets (either LLMgCode or HaCode). The prompt is shown in lines 2-3, and the code snippets to be calculated are in line 6. Based on the prompt, LLM calculates the probability of “  ” (Tab) as the first token is 0.20. Based on the prompt and “  ”, the LLM calculates the probability of “return” as the second token is 0.31. Followed by “ tf” and “.train” are 0.98 and 0.99, respectively. The higher the probability of a token, the more confident the model is about generating this token. Therefore, the perplexity of the model in understanding this token is low. Perplexity is a probability-based measure and has many implementations (e.g., log of probability (Solaiman et al., 2019), log of rank (Ippolito et al., 2019), and log of probability divided by log of rank (Su et al., 2023)) in which the log of probability is common. For example, the perplexity of CodeLlama for “return” is $ln(0.31)=-1.18$. The perplexity of a model for a piece of code is usually the average perplexity of tokens.
@@
When the code is natural and coherent, the model is confident in predicting the next token based on the prompts and the existing tokens, resulting in a lower perplexity. According to the mechanism of LLMs for code generation, LLMs generate code in the unit of a token through understanding previous tokens, calculating the probability of each possible next token, and applying the sampling strategy to select one with high probability as the output. Compared with LLMgCode, HaCode is more casual and personalized, making it more unpredictable. Therefore, it is reasonable that LLMgCode has lower perplexity than HaCode. Due to its explainability and scalability (e.g., log of probability, log of rank), Perplexity has been widely used (Solaiman et al., 2019; Gehrmann et al., 2019; Ippolito et al., 2019; Mitchell et al., 2023; Su et al., 2023).
@@
Although Perplexity is becoming the most popular method, evaluating Perplexity is limited to detection accuracy. For example, Shi et al. (2025) evaluated the proposed DetectCodeGPT using AUC (Area Under the Curve) on Python only. Yet it is unclear what the detection accuracy of DetectCodeGPT is other than Python. DetectCodeGPT detects only the code generated by LLMs with fewer than 7B parameters; its effectiveness on the commonly used commercial models such as GPT-4o and Gemini-1.0 is unclear. Moreover, DetectCodeGPT uses the same model for code generation and perplexity calculation. As it is hard to know in advance what model is for code generation, the effectiveness of DetectCodeGPT in a wide range of contexts is unclear. The limited evaluation of Perplexity does not reflect its effects well in practice.
@@
Toward realistic evaluations of Perplexity, we conduct a large-scale investigation on Perplexity from three perspectives: detection accuracy, detection speed, and generalization capability. Detection accuracy is the most fundamental goal of detecting LLMgCode, while it might not be good between situations, e.g., an effective method for C++ may not suit Python. In this article, we evaluate the accuracy in various situations from three aspects. Specifically, the detection accuracy (effectiveness) is evaluated from the aspects of the programming language (C, C#, C++, Go, Java, Python, Ruby, and Rust), degree of difficulty (easy, middle, and hard), and scale of solution (small, middle, and large). Efficiency greatly affects the user experience and large-scale applications; thereby, the detection speed (seconds/sample) of Perplexity should be considered. Since new powerful LLMs are emerging, developers utilize various LLMs to assist in programming, the generalization capability of Perplexity is important. The generalization capability is evaluated from the aspect of code generation models (GPT-3.5, GPT-4o, Gemini-1.0, and Llama-3.1). We hope the multi-perspective efficacy evaluations can reflect the efficacy of Perplexity in practice. With experimental evaluation, we first devised a dataset containing 729 programming puzzles, each accompanied by at least two HaCode solutions in each of the eight programming languages available in CodeNet (Puri et al., 2021). To obtain LLMgCode solutions, we used GPT-3.5 to generate two solutions per puzzle in each language, resulting in a total of 11,664 LLMgCode solutions. For each specific evaluation aspect, we randomly sampled 500 LLMgCode and 500 HaCode solutions for each category. For example, when it comes to programming languages, we sampled 500 LLMgCode and 500 HaCode solutions for each language. To evaluate the generalization capability of Perplexity, we generated 500 LLMgCode solutions using each of GPT-4o, Gemini-1.0, and Llama-3.1, ensuring a consistent sample size with other evaluation aspects. Specifically, we randomly selected 32 puzzles from the dataset and used each LLM (GPT-4o, Gemini-1.0, and Llama-3.1) to generate two solutions per puzzle in each language. Each LLM thus produced 512 LLMgCode solutions, from which we randomly sampled 500 to form the final set of its solutions. Then, we compared eight perplexity-based methods against two feature-based methods and one pre-training-based method that have been used to detect LLMgCode from HaCode.
@@
The experimental results show that (1) for detection accuracy, the best-performing Perplexity performs well in detecting LLMgCode in C/C++ (AUC=91%) and large-scale code (AUC=88%) but still does not perform as well as the other two methods regarding programming language, degree of difficulty, and scale of solution; (2) for detection speed, Perplexity without perturbation is as efficient as others, while Perplexity with perturbation have a relatively slow efficiency, with detection speed ranging from 10 to 150 seconds per sample; (3) for generalization capability, Perplexity perform better than others across different LLMs, with AUC ranging from 80% to 82%. Based on experimental results, we provide implications of Perplexity from both positive and negative perspectives. On the negative side, Perplexity is unsuitable for high-level programming languages, short code snippets, and just-in-time detection. On the positive side, Perplexity exhibits good generalization capability, which performs well for C/C++, and has good interpretability. Finally, we provide three directions for improving Perplexity in the future and recommendations for the application of Perplexity.
@@
  • Experimental design: An experimental design to investigate the detection accuracy, detection speed, and generation capability of Perplexity in detecting LLMgCode.
  • Findings: A wide range of findings that illustrate the advantages and limitations of Perplexity in detecting LLMgCode.
@@
The remainder of this article is organized as follows. Section 2 presents background and related work. Section 3 illustrates motivating examples. Section 4 and Section 5 elaborate on the research methodology and results, respectively. Section 6 discusses implications. Finally, we present threats to validity in Section 7 and conclude this article in Section 8.
2. Background and Related Work
@@
This section presents an overview of LLMs for code generation and reviews studies that apply Perplexity for code authorship attribution.
2.1. Background: Large Language Models for Code Generation
@@
Since LLMs were proposed, they have been widely applied in many tasks across a variety of domains (Bommasani et al., 2021; Wei et al., 2022; Zhao et al., 2023; Wu et al., 2023). One of the most exciting applications of LLMs is code generation (Brown et al., 2020; Dong et al., 2024; Hong et al., 2024; Zhang et al., 2023; Zhou et al., 2023). LLMs are capable of generating source code based on prompts. The prompts consist of natural and programmatic descriptions of code functions, retractions, input and output formats, or any other elements. Following the prompts, an LLM generates tokens sequentially until it arrives at a stop indicator or the token limit.
@@
Many studies show that LLMgCode helps developers save time (Coignion et al., 2024), reduce effort (Ziegler et al., 2022), and improve productivity (Siddiq et al., 2024) in software development. Moreover, LLMgCode help ease developers as they sometimes suffer from heavy workload (Dunne et al., 2024). With LLMgCode, developers can pay more attention to important and urgent tasks. Because of these benefits, LLMgCode has been widely used and has even become a must-have in modern software development (Spiess et al., 2025; North et al., 2024). A survey made with 500 USA developers showed that 92% of respondents use LLMgCode for official assignments and personal use (Shani and Staff, 2023).
@@
Although LLMs bring many benefits, the quality and security of LLMgCode remain challenging (Pearce et al., 2022; Yu et al., 2024b; Dakhel et al., 2023). One of the causes is that the corpus used for training LLMs is not qualified and secure enough (Asare et al., 2023). Therefore, LLMgCode might be buggy and vulnerable (Schuster et al., 2021; Jesse et al., 2023). For example, Fu et al. (2022) assessed the security of LLMgCode in six common programming languages and found that 156 out of 435 code segments contained security vulnerabilities. The proportion of vulnerabilities in C++ and Go code accounts for 46.1% and 45.0%, respectively. Mastropaolo et al. (2023) assessed the robustness of LLMgCode and found that slight modifications to the natural description while still maintaining semantic equivalence would result in LLMs generating different results, occurring in 46% (408/892) of the total samples. Jimenez et al. (2023) assessed the capability of large language models to solve 2294 real-world problems (Issues) on GitHub and found that the pass rate of code generated using GPT-4 has the lowest pass rate–0%, using Claude 2 has the highest pass rate–1.96%. In addition, LLMgCode might be similar or even unchanged from existing code snippets in the training corpus (Guo et al., 2024). As the training corpus might include a significant amount of restrictive code (Yu et al., 2023), there are copyright infringement risks for LLMgCode.
2.2. Related Work: Perplexity for Code Authorship Attribution